OpenStudy (anonymous):

Line L passes through the points (1, 5) and (5, 8). What is the slope of a line that is perpendicular to line L? -3/4 -4/3 3/4 4/3

hartnn (hartnn):

can you find the slope of a line given 2 points??

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

m=3/4 so thats my final answer?

hartnn (hartnn):

thats the slope of line L but we need the slope of a line that is perpendicular to line L!

hartnn (hartnn):

the slope of perpendicular lines have product = -1

hartnn (hartnn):

so m*(3/4) = -1 find m

OpenStudy (anonymous):

m=-4/3

hartnn (hartnn):

thats correct! :)

OpenStudy (anonymous):

Thank u!

hartnn (hartnn):

welcome ^_^